OpenStudy (anonymous):

Help!! The Proton has a mass of 1.67 x 10^-24g and a diameter of 1.000 x 10^-13. Calculate the density of the nucleus assuming it so be spherical in shape.

OpenStudy (anonymous):

density = Mass/volume Volume of Sphere is \[4/3pir ^{2}\] r= d/2

OpenStudy (anonymous):

where did the 4 come from?

OpenStudy (anonymous):

its formula of Volume of sphere = (4/3) [pi] r3

OpenStudy (anonymous):

ops the above one is cube ,not square \[4/3pir ^{3}\]

OpenStudy (anonymous):

Now just put the values ^_^